OpenStudy (cypress7):

Express answer in lowest form -1/3 + W = 5 3/8 W=

OpenStudy (jhannybean):

Is the question written out correctly? Hoow can they equal eachother?

OpenStudy (cypress7):

oh no sorry I meant -1/3 + w

OpenStudy (jhannybean):

\[-\frac{1}{3} + w=5 \frac{3}{8}\] Is that your question?

OpenStudy (cypress7):

Yes

OpenStudy (jhannybean):

To change the complex fraction, we multiply 8 x 5 and add in 3.\[\frac{(8\cdot 5)+3}{8}\]

OpenStudy (jhannybean):

What does that give you?

OpenStudy (cypress7):

43 over 8

OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

\[-\frac{1}{3} + w= \frac{43}{8}\] Now you must add \(+\dfrac{1}{3}\) to both sides.

OpenStudy (jhannybean):

So you will get \[w=\frac{43}{8}+\frac{1}{3}\]\[w=\frac{(43\cdot 3) +(8\cdot 1)}{24}\] Now just simplify.

OpenStudy (cypress7):

Thank you

OpenStudy (jhannybean):

No problem :)